# PCL-Indexability and Whittle Index for Restless Bandits with General Observation Models

### v3

@@ 2.5 Partial Conservation Law
(ii) $\forall\boldsymbol{\omega_{i}},\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}}),\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$, there must be $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$ or $\Omega_{\boldsymbol{\omega_{j}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}^{C}$, and if $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, then $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}^{C}$.
Furthermore, the entire full subset chain of $\tilde{\Omega}$ is denoted as $\mathcal{C}(\tilde{\Omega})$.
Now assume $\tilde{\Omega}$ satisfies $\mathcal{C}(\tilde{\Omega})\neq\emptyset$. For $C\in\mathcal{C}(\tilde{\Omega})$, we can use the equation (33) to formally construct variable $\lambda_{\boldsymbol{\omega_{i}}}^{C}=\frac{W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{C}}}{A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}^{C}}}$. With the above concept, we provide the definition of PCL-indexability for countable state problems as follows:
@@
We call the problem is PCL-indexable w.r.t. $\tilde{\Omega}$ if
(i) The problem satisfies PCL w.r.t. $\tilde{\Omega}$;
(ii) $\exists!C\in\mathcal{C}(\tilde{\Omega})$, s.t.$\forall\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$ satisfying $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, there is $\lambda_{\boldsymbol{\omega_{i}}}^{C}<\lambda_{\boldsymbol{\omega_{j}}}^{C}$.
Condition (i) in Definition 6 characterizes part of the boundary of the achievable region through the constraints between performance measures for an infinite dimensional linear programming problem, and condition (ii) means that the objective function can take its maximum value at a unique vertex in the achievable region while the policy corresponding to this vertex is a state priority policy obtained by sorting according to the size of the index $\lambda_{\boldsymbol{\omega}}$. Next, we show that if the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then we can construct a unique full subset chain of $\tilde{\Omega}_{0}:=\tilde{\Omega}\cup\{\{0\}\cup\Omega,\Omega\in\tilde{\Omega}\}$. Assume $\exists!C\in\mathcal{C}(\tilde{\Omega})$, s.t.$\forall\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$ with $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, we have $\lambda_{\boldsymbol{\omega_{i}}}^{C}<\lambda_{\boldsymbol{\omega_{j}}}^{C}$. We can construct the full subset chain $C^{\prime}$ of $\tilde{\Omega}_{0}$ as follows:
@@
Assume the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then for the unique full subset chain $C^{\prime}$ constructed above, it has the following index characterization:
@@
If $\Omega_{\boldsymbol{\omega_{i}}}$ has the form of (45), then it clearly satisfies the definition of the full subset chain. We will prove that the full subset chain that satisfies condition (ii) in Definition 6 is exactly in the form of (45). Assume that there exists $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ such that $\lambda_{\boldsymbol{\omega_{j}}}>\lambda_{\boldsymbol{\omega_{i}}}$, so that $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{i}}}$. Since $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{j}}}$, by condition (ii) in Definition 5, $\Omega_{\boldsymbol{\omega_{j}}}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}$. By condition (ii) in Definition 6, we have $\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}$, this contradicts with $\lambda_{\boldsymbol{\omega_{j}}}>\lambda_{\boldsymbol{\omega_{i}}}$. Similarly, assume that there exists $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ such that $\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}$, but $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}$, then $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{j}}}$ implies that $\Omega_{\boldsymbol{\omega_{i}}}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}$, which contradicts with the hypothesis. $\Box$
@@
$(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}})\frac{W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}}{A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}}$.
@@
$W_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}(A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}})=A_{\boldsymbol{\omega_{i}}}^{\Omega_{\boldsymbol{\omega_{i}}}}(W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}-W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash\{\boldsymbol{\omega_{i}}\}}).$
If $\boldsymbol{\omega_{j}}=0$, then we have $A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}=A_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash{\{\boldsymbol{\omega_{i}}\}}}=1$ and $W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{\boldsymbol{\omega_{i}}}\backslash{\{\boldsymbol{\omega_{i}}\}}}=\lambda$; if $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ and $\boldsymbol{\omega_{i}}=0$, then by definition, $A_{\boldsymbol{\omega_{j}}}^{\Omega_{0}}=A_{\boldsymbol{\omega_{j}}}^{\Omega_{0}\backslash{\{0\}}}$ and $W_{\boldsymbol{\omega_{j}}}^{\Omega_{0}}=W_{\boldsymbol{\omega_{j}}}^{\Omega_{0}\backslash{\{0\}}}$. In both cases, the lemma is clearly true. $\Box$
@@ 2.6 Optimality of Index Policy
Now we address the main issue whether the sequence of $\{\lambda_{\boldsymbol{\omega}}\}$-priorities when the problem satisfies PCL-indexability achieves optimality. Since the problem satisfies PCL, the partial conservation law inequality describes a region for an infinite-dimensional linear programming problem. First, we give a brief introduction to the basic theory of infinite-dimensional linear programming. Let $X$ and $Y$ be real linear vector spaces and the corresponding positive cones are $P_{X}$ and $P_{Y}$. Let $X$ be partially ordered by relation $\geq$, defined by $x\geq y$ if $x-y\in P_{X}$. Let $X^{*}$ be the dual space of $X$. For any $x^{*}\in X^{*}$, $x^{*}$ is a linear functional on $X$ and the image of $x\in X$ is denoted by $\langle x,x^{*}\rangle$. Similarly, space $Y$ for the dual problem also derives $Y^{*}$ and $P_{Y^{*}}$. Let $A$ be a linear map from $X$ to $Y$, $y\in Y$ and $x^{*}\in X^{*}$. The primal linear problem is
@@
Similar to the case of finite dimensions, weak duality relation can be established between (LP) and (LD).
@@
(Anderson and Nash, 1987) Assume $x$ is feasible for (LP) and $y^{*}$ is feasible for (LD), then
@@
For $A:X\rightarrow Y,~{}x^{1}\mapsto Ax^{1}$, define $Ax^{1}(\Omega)=\sum_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1},\Omega\in\tilde{\Omega}$. Let $Y$ be the linear vector space spanned by $b$ and $Ax^{1},~{}x^{1}\in X$. The positive cone is $P_{Y}=\{y\in Y:y(\Omega)\geq 0,\Omega\subset\Omega(\boldsymbol{\omega_{0}}),y(\Omega(\boldsymbol{\omega_{0}}))=0\}$. By Lemma 2, $\Omega_{\boldsymbol{\omega_{i}}}=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}\leq\lambda_{\boldsymbol{\omega_{i}}}\}$. Let $\Omega_{\boldsymbol{\omega_{i}}}^{-}:=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}\}$. In infinite dimensional problems, for any $y\in Y$, define $\lambda^{*}\in Y^{*}$ as follows:
@@
Let $x^{W}\in X$ be the performance measure given by the $\{\lambda_{\boldsymbol{\omega_{i}}}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$-priority policy. Our goal is to prove that $x^{W}$ and $\lambda^{*}$ are optimal solutions to (LP) and (LD), respectively. Due to the assumption that the problem satisfies PCL-indexability w.r.t. $\tilde{\Omega}$, it satisfies the PCL condition and $x^{W}$ is a feasible solution of (LP). By the weak duality principle, we need to prove the following two statements:
@@
Sequence $\{\lambda_{\boldsymbol{\omega}}\}_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})\cup\{0\}}$ has finitely many limit points.
@@
For $S\subset\mathcal{M}$, let $R_{T,i}^{S}$ denote the expected discounted reward obtained from initial belief state $i$ under the $S$-priority policy over the first $T$ time steps. Then we have $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{S}=R_{i}^{S}$. In fact, this limit process is independent of $j$, i.e. $R_{T,i}^{S}\rightrightarrows R_{i}^{S}(T\rightarrow+\infty)$.
@@
To prove Theorem 5, we only need to consider the case where the sequence $\{\lambda_{i}\}_{i\in\mathcal{M}}$ has one limit point. For notation convenience, we renumber the states in $\mathcal{M}$ as $\{*,1,2,\cdots\}$. The state indices in the sequence have the following relationship:
[…]
In this figure, $\lambda_{*}$ is the unique limit point of $\{\lambda_{i}\}_{i\in\mathcal{M}}$ and $\{\lambda_{2n-1}\}\searrow\lambda_{*}$ while $\{\lambda_{2n}\}\nearrow\lambda_{*}$. Recall that $\Omega_{i}=\{j\in\mathcal{M}:\lambda_{j}\leq\lambda_{i}\}$, we have the following lemma:
@@
We first prove that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}}$. We will use induction to show that for any $T$, $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$. When $T=1$, if $i\notin\Omega_{*}^{c}$, then for any $n$, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=0$; If $i\in\Omega_{*}^{c}$, since $\lambda_{*}$ is the limit point, there $\exists N$ s.t. $i\in\Omega_{2n-1}^{c}$ when $n>N$. In this case, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=R_{i}$. Thus $\lim\limits_{n\rightarrow+\infty}R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}$ holds for every $i\in\mathcal{M}$.
@@
For any $\varepsilon>0$, since $p_{i*}^{0}+\sum\limits_{j=1}^{+\infty}p_{ij}^{0}=1$, there $\exists J$ s.t. $\sum\limits_{j=J+1}^{+\infty}p_{ij}^{0}\leq\frac{\varepsilon(1-\beta)}{4\beta C}$. For $j=*,1,\cdots,J$, by the induction hypothesis, $\exists N(j)$ s.t. when $n>N(j)$, $|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|\leq\frac{\varepsilon}{2\beta}$. Let $N=\max\limits_{j\in\{*,1,\cdots,J\}}N(j)$, then when $n>N$, we have
@@
(ii) If $i\in\Omega_{*}^{c}$, then there $\exists N$ s.t. $i\in\Omega_{n}^{c}$ when $n>N$. In this case, we have
@@
By Theorem 7.11 in Rudin et al. (1976), $\lim\limits_{n\rightarrow+\infty}\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=\lim\limits_{T\rightarrow+\infty}\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$. By Lemma 6,
@@
Using similar arguments, it can be shown that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n}^{c}}=R_{i}^{\Omega_{*}^{-c}},~{}\forall i\in\mathcal{M}$. $\Box$
@@
Thus we have $\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}}$. $\Box$
@@
We will separately consider three cases where $j$ is odd, $j=*$ and $j$ is even.
(i) If $j$ is odd, we can assume $j=2N-1$, then
@@
[…]
@@
It shows that $W_{j}^{\Omega_{j}}=R_{j}+\sum\limits_{i:\lambda_{i}>\lambda_{j}}(W_{j}^{\Omega_{i}}-W_{j}^{\Omega_{i}^{-}})$ and Theorem 5 is proved. $\Box$
@@
$\forall x^{1}\in X$, following from Proposition 4.3 in Frostig and Weiss (1999), we have
@@
Thus $\langle x^{W},A^{*}\lambda^{*}-R\rangle=0$Box
From the proof of Lemma 9, we know that $\lambda^{*}$ satisfies $A^{*}\lambda^{*}-\boldsymbol{R}\in P_{X^{*}}$. Therefore, the remaining condition in the weak duality theorem is $-\lambda^{*}\in P_{Y^{*}}$. To prove this statement, we need to show that, $\forall y\in P_{Y},\langle y,~{}\lambda^{*}\rangle\leq 0$. In fact,
@@
For any $\varepsilon>0$, $\sum\limits_{i\in\mathcal{M}}|x_{i}^{1}|<+\infty$ shows that $\sum\limits_{i=1}^{+\infty}|x_{2i}^{1}|<+\infty$, and $\exists$ constant $B$ s.t. $|x_{i}^{1}|\leq B$ holds for every $i\in\mathcal{M}$. Then there $\exists I$, s.t. $\sum\limits_{i=I+1}^{+\infty}|x_{2n}^{1}|<\frac{(1-\beta)\varepsilon}{4(1+\beta)}$. Since $\lim\limits_{n\rightarrow+\infty}A_{i}^{\Omega_{2n-1}}=A_{i}^{\Omega_{*}}$, for any $1\leq i\leq I$, there $\exists N(i)$ s.t. $|A_{2i}^{\Omega_{2n-1}}-A_{2i}^{\Omega_{*}}|<\frac{\varepsilon}{2(I+1)B}$ when $n>N(i)$. Similarly, $\exists N(*)$ s.t. $|A_{*}^{\Omega_{2n-1}}-A_{*}^{\Omega_{*}}|<\frac{\varepsilon}{2(I+1)B}$ when $n>N(*)$. Let $N=\max\limits_{i\in\{*,1,\cdots,I\}}N(i)$, then when $n>N$, we have

### v4

@@ 2.5 Partial Conservation Law
(ii) $\forall\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}),\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$, there must be $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$ or $\Omega_{\boldsymbol{\omega_{j}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}^{C}$, and if $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, then $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}^{C}$.
[…]
Definition 5 aims to find an inclusive but separable relation between the subsets of belief states. Separability (if $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$, then $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}^{C}$) ensures that different belief states have different Whittle indices while inclusiveness ($\forall\boldsymbol{\omega_{i}},\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}),\boldsymbol{\omega_{i}}\neq\boldsymbol{\omega_{j}}$, there must be $\Omega_{\boldsymbol{\omega_{i}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}^{C}$ or $\Omega_{\boldsymbol{\omega_{j}}}^{C}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}^{C}$) leads to a well-order by Whittle indices detailed below.
[…]
@@
We call the problem PCL-indexable w.r.t. $\tilde{\Omega}$ if
(i) The problem satisfies PCL (Definition 4) w.r.t. $\tilde{\Omega}$;
[…]
@@
Assume the problem is PCL-indexable w.r.t. $\tilde{\Omega}$, then the unique full subset chain constructed above has the following index characterization:
@@
If $\Omega_{\boldsymbol{\omega_{i}}}$ has the form of (45), then it clearly satisfies the definition of the full subset chain. We will prove that the full subset chain that satisfies condition (ii) in Definition 6 is exactly in the form of (45). Assume that there exists $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ such that $\lambda_{\boldsymbol{\omega_{j}}}>\lambda_{\boldsymbol{\omega_{i}}}$, so that $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{i}}}$. Since $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{j}}}$, by condition (ii) in Definition 5, $\Omega_{\boldsymbol{\omega_{j}}}\subsetneq\Omega_{\boldsymbol{\omega_{i}}}$. By condition (ii) in Definition 6, we have $\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}$, this contradicts with $\lambda_{\boldsymbol{\omega_{j}}}>\lambda_{\boldsymbol{\omega_{i}}}$. Similarly, assume that there exists $\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}})$ such that $\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}$, but $\boldsymbol{\omega_{j}}\notin\Omega_{\boldsymbol{\omega_{i}}}$, then $\boldsymbol{\omega_{j}}\in\Omega_{\boldsymbol{\omega_{j}}}$ implies that $\Omega_{\boldsymbol{\omega_{i}}}\subsetneq\Omega_{\boldsymbol{\omega_{j}}}$, which contradicts with the hypothesis. ∎
Intuitively, Lemma 2 is just a formal statement that the full subset chain defined in Definition 5 with a well-ordered index function defined in Definition 6 leads to the construction of a series of monotonically increasing state subsets such that each state has a well-defined priority index.
@@
[…]
@@
[…]
@@ 2.6 Optimality of Index Policy
Now we address the main issue whether the sequence of $\{\lambda_{\boldsymbol{\omega}}\}$-priorities when the problem satisfies PCL-indexability achieves optimality. Since the problem satisfies PCL, the partial conservation law inequality describes a region for an infinite-dimensional linear programming problem. First, we give a brief introduction to the basic theory of infinite-dimensional linear programming. Let $X$ and $Y$ be real linear vector spaces and the corresponding positive cones are $P_{X}$ and $P_{Y}$. Let $X$ be partially ordered by relation $\geq$, defined by $x\geq y$ if $x-y\in P_{X}$. Let $X^{*}$ be the dual space of $X$. For any $^{*}\in X^{*}, xx^{*}$ is a linear functional on $X$. Define operator $\langle x,x^{*}\rangle$ to be the image of $x\in X$ under $x^{*}\in X^{*}$. Similarly, space $Y$ for the dual problem also induces $Y^{*}$ and $P_{Y^{*}}$. Let $A$ be a linear map from $X$ to $Y$, $y\in Y$ and $x^{*}\in X^{*}$. The primal linear problem is
@@
Similar to the case of finite dimensions, a weak duality relation can be established between (LP) and (LD).
@@
(anderson1987linear) Assume $x$ is feasible for (LP) and $y^{*}$ is feasible for (LD), then
@@
For $A:X\rightarrow Y$, define $Ax^{1}(\Omega)=\sum_{\boldsymbol{\omega}\in\Omega}A_{\boldsymbol{\omega}}^{\Omega}x_{\boldsymbol{\omega}}^{1},\Omega\in\tilde{\Omega}$. Let $Y$ be the linear vector space spanned by $b$ and $Ax^{1},~x^{1}\in X$. The positive cone is $P_{Y}=\{y\in Y:y(\Omega)\geq 0,\Omega\subset\Omega(\boldsymbol{\omega_{0}}),y(\Omega(\boldsymbol{\omega_{0}}))=0\}$. By Lemma 2, $\Omega_{\boldsymbol{\omega_{i}}}=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}\leq\lambda_{\boldsymbol{\omega_{i}}}\}$. Let $\Omega_{\boldsymbol{\omega_{i}}}^{-}:=\{\boldsymbol{\omega_{j}}\in\Omega(\boldsymbol{\omega_{0}}):\lambda_{\boldsymbol{\omega_{j}}}<\lambda_{\boldsymbol{\omega_{i}}}\}$. In infinite dimensional problems, for any $y\in Y$, define $\lambda^{*}\in Y^{*}$ as follows:
@@
The above $\lambda^{*}$ is well defined since the right side of the equation is a linear functional on $Y$ and two different linear functionals differ at least at one point. Let $x^{W}\in X$ be the performance measure given by the $\{\lambda_{\boldsymbol{\omega_{i}}}\}_{\boldsymbol{\omega_{i}}\in\Omega(\boldsymbol{\omega_{0}})}$-priority policy. Our goal is to prove that $x^{W}$ and $\lambda^{*}$ are optimal solutions to (LP) and (LD), respectively. Due to the assumption that the problem satisfies PCL-indexability w.r.t. $\tilde{\Omega}$, it satisfies the PCL condition and $x^{W}$ is a feasible solution of (LP). By the weak duality principle, we need to prove the following two statements:
@@
The sequence $\{\lambda_{\boldsymbol{\omega}}\}_{\boldsymbol{\omega}\in\Omega(\boldsymbol{\omega_{0}})\cup\{0\}}$ has finitely many limit points.
@@
[…]
@@
To prove Theorem 5, we only need to consider the case where the sequence $\{\lambda_{i}\}_{i\in\mathcal{M}}$ has one limit point. For notational convenience, we renumber the states in $\mathcal{M}$ as $\{*,1,2,\cdots\}$. The state indices in the sequence have the relationship as shown in Fig. 1.
Figure 1: Limit point
$\lambda_{2}$$\lambda_{4}\cdots$$\lambda_{*}\cdots$$\lambda_{3}$$\lambda_{1}$
In Fig. 1, $\lambda_{*}$ is the unique limit point of $\{\lambda_{i}\}_{i\in\mathcal{M}}$ and $\{\lambda_{2n-1}\}\searrow\lambda_{*}$ while $\{\lambda_{2n}\}\nearrow\lambda_{*}$. Recall that $\Omega_{i}=\{j\in\mathcal{M}:\lambda_{j}\leq\lambda_{i}\}$, we have the following lemma:
@@
We first prove that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n-1}^{c}}=R_{i}^{\Omega_{*}^{c}}$. We will use induction to show that for any $T$, $\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=R_{T,i}^{\Omega_{*}^{c}}$. When $T=1$, if $i\notin\Omega_{*}^{c}$, then for any $n$, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=0$; If $i\in\Omega_{*}^{c}$, since $\lambda_{*}$ is the limit point, $\exists N$ s.t. $i\in\Omega_{2n-1}^{c}$ when $n>N$. In this case, $R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}=R_{i}$. Thus $\lim\limits_{n\rightarrow+\infty}R_{1,i}^{\Omega_{2n-1}^{c}}=R_{1,i}^{\Omega_{*}^{c}}$ holds for every $i\in\mathcal{M}$.
@@
For any $\varepsilon>0$, since $p_{i*}^{0}+\sum\limits_{j=1}^{+\infty}p_{ij}^{0}=1$, $\exists J$ s.t. $\sum\limits_{j=J+1}^{+\infty}p_{ij}^{0}\leq\frac{\varepsilon(1-\beta)}{4\beta C}$. For $j=*,1,\cdots,J$, by the induction hypothesis, $\exists N(j)$ s.t. when $n>N(j)$, $|R_{T,j}^{\Omega_{2n-1}^{c}}-R_{T,j}^{\Omega_{*}^{c}}|\leq\frac{\varepsilon}{2\beta}$. Let $N=\max\limits_{j\in\{*,1,\cdots,J\}}N(j)$, then when $n>N$, we have
@@
(ii) If $i\in\Omega_{*}^{c}$, then $\exists N$ s.t. $i\in\Omega_{n}^{c}$ when $n>N$. In this case, we have
@@
By Theorem 7.11 in rudin1976principles, $\lim\limits_{n\rightarrow+\infty}\lim\limits_{T\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}=\lim\limits_{T\rightarrow+\infty}\lim\limits_{n\rightarrow+\infty}R_{T,i}^{\Omega_{2n-1}^{c}}$. By Lemma 6,
@@
Using similar arguments, it can be shown that $\lim\limits_{n\rightarrow+\infty}R_{i}^{\Omega_{2n}^{c}}=R_{i}^{\Omega_{*}^{-c}},~\forall i\in\mathcal{M}$. ∎
@@
Thus we have $\lim\limits_{n\rightarrow+\infty}W_{i}^{\Omega_{2n+1}}=W_{i}^{\Omega_{*}}$. ∎
@@
We will separately consider three cases where $j$ is odd, j=j=* and $j$ is even.
(i) If $j$ is odd, we can assume j=2N1j=2N-1, then
@@
$\lambda_{2N}$$\lambda_{2N+2}\cdots$$\lambda_{*}\cdots$$\lambda_{3}$$\lambda_{1}$
@@
It shows that $W_{j}^{\Omega_{j}}=R_{j}+\sum\limits_{i:\lambda_{i}>\lambda_{j}}(W_{j}^{\Omega_{i}}-W_{j}^{\Omega_{i}^{-}})$ and Theorem 5 is proved. ∎
@@
$\forall x^{1}\in X$, following from Proposition 4.3 in frostig1999four, we have
@@
Thus $\langle x^{W},A^{*}\lambda^{*}-R\rangle=0$
From the proof of Lemma 9, we know that $\lambda^{*}$ satisfies $A^{*}\lambda^{*}-\boldsymbol{R}\in P_{X^{*}}$. Therefore, the remaining condition in the weak duality theorem is $\lambda^{*}\in P_{Y^{*}}$. To prove this statement, we need to show that, $\forall y\in P_{Y},\langle y,~\lambda^{*}\rangle\leq 0$. In fact,
@@
For any $\varepsilon>0$, $\sum\limits_{i\in\mathcal{M}}|x_{i}^{1}|<+\infty$ shows that $\sum\limits_{i=1}^{+\infty}|x_{2i}^{1}|<+\infty$, and $\exists$ constant $B$ s.t. $|x_{i}^{1}|\leq B$ holds for every $i\in\mathcal{M}$. Then $\exists I$, s.t. $\sum\limits_{i=I+1}^{+\infty}|x_{2n}^{1}|<\frac{(1-\beta)\varepsilon}{4(1+\beta)}$. Since $\lim\limits_{n\rightarrow+\infty}A_{i}^{\Omega_{2n-1}}=A_{i}^{\Omega_{*}}$, for any $1\leq i\leq I$, $\exists N(i)$ s.t. $|A_{2i}^{\Omega_{2n-1}}-A_{2i}^{\Omega_{*}}|<\frac{\varepsilon}{2(I+1)B}$ when $n>N(i)$. Similarly, $\exists N(*)$ s.t. $|A_{*}^{\Omega_{2n-1}}-A_{*}^{\Omega_{*}}|<\frac{\varepsilon}{2(I+1)B}$ when n>N()n>N(*). Let $N=\max\limits_{i\in\{*,1,\cdots,I\}}N(i)$, then when $n>N$, we have